# Optimal Treatment Allocation Strategies for A/B Testing in Partially Observable Time Series Experiments

### v2

@@ 4.2 A Reinforcement Learning Approach to Learning Optimal $q$-dependent Policies
We defer the proof of Theorem 3 to Appendix A and make a few remarks: (i) Corollary 1 in Section 3.3 proves the optimality of the balanced design for AR processes. Theorem 3 extends this to (V)ARMA processes, allowing residuals to be correlated over time. (ii) The determinism, stationarity, and $q$-dependency conditions significantly reduce the search space from over $2^{T}$ to less than $2^{q+1}$, simplifying the learning of $\pi^{*}$. These conditions enable us to focus on this restricted class to find $\pi^{*}$ by minimizing (4.3). (iii) The proof of Theorem 3 draws from existing proofs establishing the Markov and stationarity properties of the optimal policy in RL (see, e.g., Puterman,, 2014; Ljungqvist and Sargent,, 2018). A crucial step in our proof is to construct an MDP and establish the equivalence between learning the optimal policy that maximizes the average reward in this MDP and identifying the optimal treatment allocation strategies that minimize (4.3). To elaborate, we introduce the following sequence of state-action-reward triplets $(\mathbf{S}_{t},A_{t},R_{t})_{t>q}$:
@@ 5 Experiments
We demonstrate the finite sample performance of our proposed methods using a dispatch simulator (Xu et al.,, 2018) and two city-scale real datasets from a ride-sharing company. Our objectives are to (i) validate the effectiveness of the proposed efficiency indicators in comparing the commonly used AD, UR, and AT designs and (ii) conduct comparisons among various types of designs:
@@
  • The commonly used AD, UR, and AT treatment allocation strategies;
@@
We note that Greedy is commonly used in online RL for regret minimization. TMDP and NMDP are variants of AD designs that are proven to be optimal under these respective model assumptions. Finally, Switch is a variant of AT design that switches back and forth over a fixed period rather than at every decision point. The optimal duration of each switch is determined by the order of the carryover effect, and we select the best duration from $\{2,5,10\}$ to report.
5.1 Synthesis Data from a Dispatch Simulator
Environment. We simulate a small-scale synthetic ride-sharing environment as in Xu et al., (2018) and Li et al., (2023), where drivers and customers interact within a 9$\times$9 spatial grid over 20 time steps per day:
@@ 5.1 Synthesis Data from a Dispatch Simulator
  • Policies. We compare two order dispatching policies: (i) a conventional distance-based policy that matches idle drivers with unassigned orders by minimizing their total distances at each time, and (ii) an MDP-based policy (Xu et al.,, 2018) that solves the matching problem by maximizing the long-term benefits of the ride-sharing platform rather than focusing on total distances at each current time.
@@
(b) MSE of Sequential Designs
Figure 4: (a) Scatter plots of ATE estimators against the two efficiency indicators $\text{EI}_{\text{AD}}$ and $\text{EI}_{\text{AT}}$. (b) Violin plots of the MSEs of ATE estimators under different treatment allocation strategies in the synthesis example. The designs are ranked in descending order from left to right regarding average MSEs. The positions of the middle points in each violin denote the mean, while the black solid lines indicate the standard deviation.
@@
Table 1: Average MSE over Different Treatment Allocation Strategies on Synthesis Data.
Implementation. The outcome of interest is set to the revenue earned at each time step. In addition to this outcome, we include two other variables in the observation: the number of unassigned orders and the number of idle drivers each time. Implementing both the proposed efficiency indicators and designs requires estimating the AR and MA parameters. To this end, we first generate a historical dataset that lasts for 50 days. Next, we apply the VARMA model to fit this dataset to estimate the AR and MA parameters. The optimal AR and MA orders, $p^{*}$ and $q^{*}$, are selected using AIC, resulting in $p^{*}=q^{*}=2$. Using these estimators, we compute the proposed efficiency indicators and proceed to implement the proposed designs, comparing them against other previously mentioned designs. Specifically, for each design, we generate 50 days of experimental data and apply the controlled VARMA model to this dataset to estimate the ATE. Finally, we repeat the entire procedure 30 times to compute the MSE of the ATE estimator under each design. The oracle ATE is evaluated via the Monte Carlo method, resulting in a value of 2.24, leading to a 6% improvement.
@@
Results. We visualize the efficiency indicators across 50 simulation replications and the MSEs of ATE estimators under different designs in Figure 4. The values of these MSEs are detailed in Table 1. The results are summarized as follows:
@@
  • Efficiency Indicators. As shown in Figure 4 (a), the efficiency indicators for the AD design ($\text{EI}_{\text{AD}}$) are all negative, while those for the AT design ($\text{EI}_{\text{AT}}$) are mostly positive. According to Corollary 3, this pattern suggests that the AD design is likely more efficient than AT and UR in this simulation environment. The MSEs in Figure 4 (b) and Table 1 substantiate that both AT and UR designs result in significantly higher MSEs in estimating the ATE compared to AD. These findings highlight the effectiveness of the proposed efficiency indicators when comparing the three designs.
@@
  • Designs. As seen in Figure 4 (b) and Table 1, our proposed CO and RL designs lead to the most efficient ATE estimators in this synthetic example. When solving (4.2) to implement the CO, the argmin $\alpha^{*}$ is found to be exactly 1, indicating that CO identifies the AD design. The proposed RL design closely mirrors the CO design as well. TMDP and NMDP outperform the commonly used AT, UR, AD, Greedy, and Switch but are inferior to our proposed designs. Although Greedy is effective in online experiments for regret minimization by balancing the exploration-exploitation trade-off, it does not necessarily optimize the performance of the resulting ATE estimator. Finally, while Switch is competitive, it is less efficient than our proposed optimal designs.
Figure 5: Trend of State Observations: driver online time, number of calls, and reward (the drivers’ income) on City 1 (the first row) and City 2 (the second row) across 240 hours (10 days).
@@ 5.2 Real Data-based Analyses
Data. We use two real datasets from two different cities, provided by a world-leading ride-sharing company, to create simulation environments for investigating the finite sample performance of the proposed efficiency indicators and designs. We do not disclose the names of the cities or the company for privacy concerns. Both datasets are generated under A/A experiments, where a single order dispatching strategy is consistently deployed over time. Each dataset contains 40 days of data and is summarized as a three-dimensional time series. The first dimension records the drivers’ total income at each time interval, serving as the outcome. The last two elements are the number of order requests and drivers’ online time at each time interval, respectively, measuring the demand and supply of the market. The time units in the datasets differ, with the first being 30 minutes and the second being one hour. See Figure 5 to visualize these three-dimensional time series.
Bootstrap-based Simulation. Figure 5 reveals clear daily trends in both time series, with a significant rise and a subsequent decline in driver income and the number of call orders during the morning and evening peak hours. To effectively capture these seasonal patterns, we incorporate a time dummy variable $D_{t}$ into our controlled VARMA model to fit the three-dimensional observation. This variable is set to one during peak hours between 8 am to 8 pm and zero otherwise.
Next, we employ the parametric bootstrap to create synthetic data. Specifically, we first fit the following VARMA-X model to the three-dimensional time series $\mathbf{Y}_{t}$, with the dummy variable $D_{t}$ serving as the exogenous variable,
$\mathbf{Y}_{t}=\bm{\mu}+\sum_{j=1}^{p}\mathbf{A}_{j}\mathbf{Y}_{t-j}+\bm{\eta}D_{t}+\mathbf{Z}_{t}.$
We next record all estimated parameters, i.e., $\{\widehat{\mathbf{A}}_{i}\}_{i=1}^{p}$, $\widehat{\bm{\Sigma}}$, $\{\widehat{\mathbf{M}}_{i=1}^{q}\}$ as well as the estimated $\bm{\eta}$, which is denoted by $\widehat{\bm{\eta}}$. Finally, we simulate synthetic time series $\{\widetilde{\mathbf{Y}}_{t}\}$ according to the following equation:
$\widetilde{\mathbf{Y}}_{t}=\sum_{j=1}^{p}\widehat{\mathbf{A}}_{j}\widetilde{\mathbf{Y}}_{t-j}+b\mathbf{1}U_{t}+\widehat{\bm{\eta}}D_{t}+\widehat{\mathbf{Z}}_{t},$ (5.1)
where $\mathbf{1}$ denotes a vector of ones, $b$ is some pre-specified parameter that determines the size of the ATE, $\{U_{t}\}_{t}$ are determined by different designs, and $\{\widehat{\mathbf{Z}}_{t}\}$ follow the estimated MA process and are generated prior to $\{\widetilde{\mathbf{Y}}_{t}\}$.
@@
Table 2: Comparison of statistical efficiency over distinct designs in estimating ATE. The bold number indicates the best result, while the underlining denotes the second-best. The symbol $\downarrow$ represents an inverse indicator, meaning that a lower value signifies a more accurate design for estimating the treatment effect. Results are averaged over 10 seeds.
  • Efficiency indicators: The results from City 1 demonstrates the effectiveness of efficiency indicators for comparing AD, UR, and AT designs. Specifically, the AT design yields a more accurate ATE estimator for this city compared to AD and UR, as evidenced by both the average MSE and the performance ranking metric. This improvement is consistent with a negative $\text{EI}_{\text{AT}}$ and a positive $\text{EI}_{\text{AD}}$. In contrast, the results in City 2 are reversed, where the AD design significantly outperforms AT with a considerably smaller average MSE and a higher ranking. Meanwhile, AD generally outperforms UR. These results are consistent with a negative $\text{EI}_{\text{AD}}$ and a positive $\text{EI}_{\text{AT}}$.
  • Designs: The violin plots in Figure 6 visualize the distribution of MSEs of ATE estimators under various treatment allocation strategies, where the width of the violin indicates the density of data points at different MSE values. The treatment allocation strategies are arranged in descending order from left to right according to the average MSE across the range of $b$. In both cities, the distributions of MSEs under the proposed CO and RL designs are more tightly centered around zero when compared to other designs. Table 2 also suggests a consistent improvement of statistical efficiency for our proposed optimal designs over alternative methods. It is also worth mentioning that the AT design achieves a competitive second-best performance ranking in City 1. In City 2, the TMDP design outperforms ours in terms of performance ranking, partly because it additionally leverages observational data to determine optimal treatments in online experiments, whereas our designs are observation-agnostic. However, neither AT nor TMDP performs well in the other city. On the contrary, the performance of our designs is more consistent and robust across different cities.
@@ 6 Discussion
In this paper, we study experimental designs for A/B testing in partially observable time series experiments. Specifically, we propose the controlled (V)ARMA model—a rich subclass of POMDPs—for fitting experimental data. We establish asymptotic MSEs of ATE estimators, derive two efficiency indicators to assess the statistical efficiency of three commonly used designs, and develop two data-driven algorithms to learn the optimal observation-agnostic design. Our work bridges several vital research areas, including time series analysis, experimental design, causal inference, RL, and A/B testing, opening up many exciting research pathways across these domains for future investigations.

### v4

@@ 4.2 A Reinforcement Learning Approach
We defer the proof of Theorem 3 to Appendix A and make a few remarks: (i) Corollary 1 in Section 3.3 proves the optimality of balanced designs for AR processes. Theorem 3 extends this to (V)ARMA processes, allowing residuals to be correlated over time. (ii) The determinism, stationarity, and $q$-dependency conditions significantly reduce the search space from over $2^{T}$ to less than $2^{q+1}$, simplifying the learning of $\pi^{*}$. These conditions enable us to focus on this restricted class to find $\pi^{*}$ by minimizing (4.4). (iii) The proof of Theorem 3 draws from existing proofs establishing the Markov and stationarity properties of the optimal policy in RL (see, e.g., Puterman,, 2014; Ljungqvist and Sargent,, 2018). A crucial step in our proof is to construct an MDP and establish the equivalence between learning the optimal policy that maximizes the average reward in this MDP and identifying the optimal treatment allocation strategies that minimize (4.4). To elaborate, we introduce the following sequence of state-action-reward triplets $(\mathbf{S}_{t},A_{t},R_{t})_{t>q}$:
@@ 5 Experiments
We demonstrate the finite sample performance of our proposed methods using two dispatch simulators (Xu et al.,, 2018; Tang et al.,, 2019) and two real datasets from a ride-sharing company. Importantly, the two simulators used in Sections 5.1 and 5.2 are based on physical models that simulate the behaviors of drivers and passengers. Therefore, the data generated from these environments does not necessarily follow the proposed controlled (V)ARMA model, providing a robust evaluation of our proposal under model misspecification.
Our objectives are to (i) validate the effectiveness of the proposed efficiency indicators in comparing AD, UR and AT; (ii) conduct comparisons among the following designs:
@@
  • The commonly used AD, UR, and AT designs;
@@
We note that Greedy is commonly used in online RL for regret minimization. TMDP and NMDP are variants of AD designs that are proven to be optimal under their model assumptions. Finally, Switch is a variant of AT design that switches back and forth over a fixed period rather than at every decision point. The optimal duration of each switch is determined by the order of the carryover effect, and we select the best duration from $\{2,5,10\}$ to report.
5.1 Synthetic Dispatch Simulator
Environment. We simulate a synthetic ride-sharing environment as in Xu et al., (2018) and Li et al., (2023), where drivers and customers interact within a 9$\times$9 spatial grid over 20 time steps per day:
@@ 5.1 Synthetic Dispatch Simulator
  • Policies. We compare two order dispatching policies: (i) a conventional distance-based policy that matches idle drivers with unassigned orders by minimizing their total distances at each time, and (ii) an MDP-based policy that solves the matching problem by maximizing the long-term benefits of the ride-sharing platform rather than focusing on total distances at each current time (Xu et al.,, 2018).
Implementation. The outcome of interest is set to the driver’s income earned at each time step. In addition to this outcome, we include two other variables in the observation: the number of unassigned orders and the number of idle drivers each time. Implementing both the proposed efficiency indicators and designs requires estimating the AR and MA parameters. To this end, we first generate a historical dataset that lasts for 50 days. Next, we apply the VARMA model to fit this dataset to estimate the AR and MA parameters. The optimal AR and MA orders, $p^{*}$ and $q^{*}$, are selected using AIC, resulting in $p^{*}=q^{*}=2$. Using these estimators, we compute the proposed efficiency indicators and proceed to implement the proposed designs, comparing them against other previously mentioned designs. Specifically, for each design, we generate 50 days of experimental data to estimate the ATE. Finally, we repeat the entire procedure 30 times to compute the MSE of the ATE estimator under each design. The oracle ATE is evaluated via the Monte Carlo method, resulting in a value of 2.24, leading to a 6% improvement.
@@
(b) MSE under different designs
Figure 4: (a) The kernel density estimation (KDE) plot for the empirical distribution of the two efficiency indicators $\text{EI}_{\text{AD}}$ and $\text{EI}_{\text{AT}}$. (b) Violin plots of MSEs of ATE estimators under different designs in environments created by the synthetic simulator. Designs are ranked in descending order from left to right according to their average MSEs in estimating the ATE. The positions of the middle points in each violin denote the average MSEs, while the lengths of violins reflect the variabilities.
@@
Table 1: Average MSEs under different designs in environments created by the synthetic simulator.
@@
Results. We visualize the efficiency indicators and the MSEs of ATE estimators under different designs in Figure 4. The values of these MSEs are detailed in Table 1. The results are summarized as follows:
@@
  • Efficiency Indicators. As shown in Figure 4(a), most of the estimated $\text{EI}_{\text{AD}}$ (colored in blue) are negative across 50 replications, while most estimated $\text{EI}_{\text{AT}}$ (in red) are positive. According to Corollary 3, this pattern suggests that the AD design is likely more efficient than AT and UR in this simulation environment. Figure 4(b) and Table 1 further verify this finding, showing that both AT and UR result in significantly higher MSEs in estimating the ATE compared to AD. These findings highlight the effectiveness of the proposed efficiency indicators when comparing the three designs.
@@
  • Designs. As seen in Figure 4(b) and Table 1, our proposed CO and RL designs lead to the most efficient ATE estimators. Meanwhile, TMDP and NMDP outperform the commonly used AT, UR, AD, Greedy, and Switch but are inferior to our proposed designs. Although Greedy is effective in online experiments for regret minimization by balancing the exploration-exploitation trade-off, it does not necessarily optimize the performance of the resulting ATE estimator.
5.2 City-level Real-data-based Dispatch Simulator
Environment. We further conduct A/B testing by using a more complicated and realistic city-level order dispatching simulator (Tang et al.,, 2019). To mimic real-world ride-sharing markets, this simulator is trained based on a historical dataset collected from a world-leading ride-sharing company in a particular city. We do not disclose the names of the cities or the company for privacy concerns. Compared with the $9\times 9$ synthetic simulator in Section 5.1, this dispatch simulator is more realistic in the following ways:
  1. 1.
    Drivers and customers interact within a real city divided into 85 hexagonal regions, as opposed to a synthetic city with grid-based rectangular regions.
  2. 2.
    Orders are generated based on historical data rather than being synthetically simulated. The order dispatching policy matches existing unassigned and new orders every 2 seconds, aligning with the company’s current practice.
  3. 3.
    Drivers are initially distributed according to their empirical distribution in the historical dataset rather than uniformly randomly distributed. Additionally, drivers assigned to orders have the option to reject them, with rejection probabilities computed by a pre-trained classification model that uses driver and order characteristics as features. Meanwhile, idle drivers may either relocate based on a random walk model trained using their historical movement data or follow the company’s instructions to move to specific locations as determined by a pre-trained repositioning algorithm. Finally, idle drivers could go offline before the next order dispatching round, while new drivers may appear online, according to historical data.
Similar to Section 5.1, the observation in this city-scale simulator is also three-dimensional, including the number of orders, the number of drivers, and the driver income, which is the outcome of interest. For each design, we conduct the online experiments over four days to estimate the ATE and replicate the experiment 30 times to calculate its root MSE.
Designs AT UR Greedy AD Switch NMDP TMDP CO RL
[…]
Table 2: Average RMSEs under different designs in environments created by the city-level real-data-based simulator.
(a) Efficiency indicators
(b) RMSE under different designs
Figure 5: (a) The KDE plot for the empirical distribution of the two efficiency indicators $\text{EI}_{\text{AD}}$ and $\text{EI}_{\text{AT}}$. (b) Violin plots of the MSEs of ATE estimators under different designs in environments created by the city-level real-data-based simulator.
Results. The empirical distribution of efficiency indicators and the root MSE (RMSE) of ATE estimators under different designs are visualized in Figure 5(a) and (b), respectively. Additionally, the values of these RMSEs are reported in Table 2 as well. We summarize the results as follows.
  • Efficiency Indicators. Figure 5(a) suggests that the estimated $\textbf{EI}_{\text{AD}}$ values are mostly negative across the 30 replications, whereas the estimated $\textbf{EI}_{\text{AT}}$ values are mostly positive. This suggests that AD is more efficient than AT and UR in this environment, which aligns with the results reported in Figure 5(b) and Table 2.
  • Designs. Figure 5 and Table 2 demonstrate the superiority of our proposed CO and RL designs, which achieves the lowest MSE among all considered designs. As mentioned earlier, both the synthetic simulator in Section 5.1 and the real-data-based simulator in this section are built based on physical models to simulate driver and customer behaviors. Even though the data from the two dispatch simulators might not follow the proposed model, our designs consistently deliver the best performance. This outperformance demonstrates the robustness of our designs against model misspecification, enabling more accurate A/B testing than existing state-of-the-art in real practice.
@@ 5.3 Real Data-based Analyses
Data. We use two real datasets from two different cities, provided by the ride-sharing company, to create simulation environments for investigating the finite sample performance of the proposed efficiency indicators and designs. Both datasets are generated under A/A experiments, where a single order dispatching strategy is consistently deployed over time. Each dataset contains 40 days of data and is summarized as a three-dimensional time series. The first dimension records the drivers’ total income at each time interval, serving as the outcome. The last two elements are the number of order requests and drivers’ online time at each time interval, respectively, measuring the demand and supply of the market. The time units in the datasets differ, with the first being 30 minutes and the second being one hour. See Figure 6 for visualizations of these three-dimensional time series.
Bootstrap-based Simulation. Figure 6 reveals clear daily trends in both time series, with a significant rise and a subsequent decline in driver income and the number of call orders during the morning and evening peak hours. To effectively capture these seasonal patterns, we incorporate a dummy variable, $D_{t}$, as an exogenous variable in our controlled VARMA model to fit the three-dimensional observation. This variable is set to one during peak hours between 8 am to 8 pm and zero otherwise.
Figure 6: Trend of observations: driver online time, number of calls, and reward (the drivers’ income) on City 1 (the first row) and City 2 (the second row) across 240 hours (10 days).
Next, we employ the parametric bootstrap to create simulated data. Specifically, we first fit the following VARMA model
$\mathbf{Y}_{t}=\bm{\mu}+\sum_{j=1}^{p}\mathbf{A}_{j}\mathbf{Y}_{t-j}+\bm{\eta}D_{t}+\mathbf{Z}_{t},$
and record all estimated parameters, i.e., $\widehat{\bm{\mu}}$, $\{\widehat{\mathbf{A}}_{i}\}_{i=1}^{p}$, $\widehat{\bm{\eta}}$, $\widehat{\bm{\Sigma}}$, $\{\widehat{\mathbf{M}}_{i=1}^{q}\}$. Next, we simulate time series $\{\widehat{\mathbf{Y}}_{t}\}$ according to the following equation:
$\widehat{\mathbf{Y}}_{t}=\widehat{\bm{\mu}}+\sum_{j=1}^{p}\widehat{\mathbf{A}}_{j}\widehat{\mathbf{Y}}_{t-j}+b\mathbf{1}U_{t}+\widehat{\bm{\eta}}D_{t}+\widehat{\mathbf{Z}}_{t},$ (5.1)
where $\mathbf{1}$ denotes a vector of ones, $b$ is some pre-specified parameter that determines the size of the ATE, $\{U_{t}\}_{t}$ are determined by different designs, and $\{\widehat{\mathbf{Z}}_{t}\}$ follow the estimated MA process and are generated prior to $\{\widehat{\mathbf{Y}}_{t}\}$.
@@
Table 3: Comparison of different designs in estimating ATE. The bold number indicates the best result, while the underlined number denotes the second-best. The symbol $\downarrow$ represents an inverse indicator, meaning that a lower value denotes a more effective design for estimating the ATE.
  • Efficiency indicators. As evidenced by both the average MSE and the performance ranking metric, the AT design yields a more accurate ATE estimator for City 1 compared to AD and UR. These findings are consistent with a negative $\text{EI}_{\text{AT}}$ and a positive $\text{EI}_{\text{AD}}$. In contrast, the results in City 2 are reversed, where the AD design significantly outperforms AT with a considerably smaller average MSE and a higher ranking. Meanwhile, AD generally outperforms UR. These results are, again, consistent with a negative $\text{EI}_{\text{AD}}$ and a positive $\text{EI}_{\text{AT}}$.
  • Designs. The violin plots in Figure 7 visualize the distribution of MSEs of ATE estimators under various designs, where the width of the violin indicates the density of data points at different MSE values. The designs are arranged in descending order from left to right according to the average MSE across the range of $b$. In both cities, the distributions of MSEs under the proposed CO and RL designs are more tightly centered around zero when compared to other designs. Table 3 also suggests a consistent improvement of statistical efficiency for our proposed optimal designs over alternatives. It is also worth mentioning that the AT design achieves a competitive second-best performance ranking in City 1. In City 2, the TMDP design outperforms ours in terms of performance ranking, partly because it additionally leverages observational data to determine optimal treatments in the online experiment, whereas our designs are observation-agnostic. However, neither AT nor TMDP performs well in the other city. On the contrary, the performance of our designs is more consistent and robust across the two cities.
@@ 6 Discussion
In this paper, we study experimental designs for A/B testing in partially observable online experiments where the data does not satisfy the Markov assumption. Specifically, we propose the controlled (V)ARMA model—a rich subclass of POMDPs—for fitting experimental data, establish asymptotic MSEs of ATE estimators, derive two efficiency indicators to assess the statistical efficiency of three commonly used designs, and develop two data-driven algorithms to learn the optimal observation-agnostic design. Our work bridges several vital research areas, including time series analysis, experimental design, causal inference, RL, and A/B testing, opening numerous exciting avenues for future research across these fields, covering theory, methodology, and applications.
6.1 Applications: Extensions to Spatially Dependent Experiments
In our ride-sharing example, we consider evaluating order dispatching policies, which are typically randomized over time and implemented in the whole city at each time, leading to temporally dependent experiments where the data generated is summarized into a single time series. However, the company is equally keen on applying different subsidizing policies in different spatial locations of the city to balance the driver supply and customer demand across the city (Shi et al., 2023a, ). These experiments are inherently spatially dependent.
Spatially dependent experiments are common in many applications that involve a group of experimental units that receive (sequential) treatments across different locations (Ugander et al.,, 2013; Baird et al.,, 2018; Johari et al.,, 2022; Leung,, 2022; Jia et al.,, 2023; Viviano et al.,, 2023; Liu et al.,, 2024; Zhan et al.,, 2024); see also Bajari et al., (2023) for a recent overview. In these experiments, in addition to the carryover effect over time, the spatial spillover effect also exists where the treatment of one experimental unit can affect the outcome of others. There is growing interest in developing causal inference methods that account for spatial interference (see Reich et al.,, 2021, for a recent overview). It would be practically interesting to integrate the proposed designs with these methodologies to adapt them to such settings.
6.2 Methodology: Model-based v.s. Model-free Approaches
Our proposal is model-based in that it employs classical time series models for estimating the ATE and designing the experiment. Alternatively, model-free methods that do not directly model the time series are equally applicable. Depending on how they estimate the ATE, these model-free methods can be roughly categorized into two types:
  1. (i)
    The first type assumes that the experimental data follows an MDP to handle carryover effects (Farias et al.,, 2022; Shi et al., 2023b, ; Cao and Zhou,, 2024) and adapts existing model-free off-policy evaluation methods developed in the RL literature for MDPs (see e.g., Liao et al.,, 2022; Kallus and Uehara,, 2022; Uehara et al.,, 2022) to evaluate the ATE.
  2. (ii)
    The second type is completely model-agnostic and employs generic importance sampling methods (see e.g., Zhang et al.,, 2013; Bojinov and Shephard,, 2019; Hu and Wager,, 2023).
Both model-based and model-free methods have their own merits. Model-based approaches are often more data efficient, leading to less variable ATE estimators. However, they can be vulnerable to model misspecification. The first type of model-free method does not rely on a specific model, but it may fail under partial observability. The second type of method allows partial observability, but it leads to more variable ATE estimators. This increased variability is undesirable for A/B testing, particularly in settings with small sample sizes and weak signals.
Our choice of the model-based approach is guided by the principle that “all models are wrong, but some are useful” (Box,, 1979). Unlike the aforementioned model-free methods, our approach not only addresses the four practical challenges mentioned in the introduction but also demonstrates its usefulness in our numerical experiments under model misspecification. Additionally, our collaborators in the ride-sharing company prefer the model-based approach for its interpretability.
Meanwhile, the proposed controlled (V)ARMA model serves as a stepping stone. It can be extended to a variety of models for future research. For instance, autoregressive fractionally integrated moving-average (ARFIMA) models could be explored to handle long-term dependencies, which are common in practice. Similarly, more general linear state space models such as those in Liang and Recht, (2023) could be considered. Despite these changes in modeling, our paper establishes a foundational framework for analysis and design. It covers both theoretical techniques, such as the small signal asymptotics, and methodological developments, including the RL algorithms, which can be adapted to these new models.
6.3 Theory: Small Signal Asymptotic Framework
At the core of our asymptotic theories is the proposed small signal asymptotic framework, which substantially simplifies the asymptotic calculations in time dependent experiments. As mentioned earlier, it aligns with our empirical observations where most improvements from new strategies are not substantial. When this assumption is violated, our designs are not guaranteed to be optimal. However, in such cases, the treatment effect becomes non-negligible. Our approach, although potentially sub-optimal, remains consistent in detecting this effect. This ensures that our design remains safe to use, regardless of whether the signal is small or large.